OpenStudy (anonymous):

HELP WITH 4 QUESTIONS PLEASE!?!?!?

OpenStudy (anonymous):

\[\frac{ 8x^3 + 6x^2 - 4x }{ -2x }\]

OpenStudy (anonymous):

\[\frac{ 6x^4 +9x^2 + 12x }{ 3x }\]

OpenStudy (anonymous):

\[\frac{ 3x^4 + 5x^2 + 2x }{ x }\]

OpenStudy (anonymous):

\[\frac{ 10x^6y^3 +20x^3y^2 }{ 5x^3y }\]